### v1

# Question-to-Knowledge: Multi-Agent Generation of Inspectable Facts for Product Mapping

Abstract
Identifying whether two product listings refer to the same Stock Keeping Unit (SKU) is a persistent challenge in e-commerce, especially when explicit identifiers are missing and product names vary widely across platforms. Rule-based heuristics and keyword similarity often misclassify products by overlooking subtle distinctions in brand, specification, or bundle configuration. To overcome these limitations, we propose Question-to-Knowledge (Q2K), a multi-agent framework that leverages Large Language Models (LLMs) for reliable SKU mapping. Q2K integrates: (1) a Reasoning Agent that generates targeted disambiguation questions, (2) a Knowledge Agent that resolves them via focused web searches, and (3) a Deduplication Agent that reuses validated reasoning traces to reduce redundancy and ensure consistency. A human-in-the-loop mechanism further refines uncertain cases. Experiments on real-world consumer goods datasets show that Q2K surpasses strong baselines, achieving higher accuracy and robustness in difficult scenarios such as bundle identification and brand–origin disambiguation. By reusing retrieved reasoning instead of issuing repeated searches, Q2K balances accuracy with efficiency, offering a scalable and interpretable solution for product integration.
Index Terms:
Large Language Models (LLMs), Multi-Agent Systems, Retrieval-Augmented Reasoning, E-commerce Data Integration, SKU Matching
I Introduction
The explosive growth of e-commerce platforms has brought renewed attention to the problem of product mapping, the task of determining whether two product listings correspond to the same Stock Keeping Unit (SKU). Accurate SKU mapping underpins critical applications such as multi-platform price monitoring, sourcing optimization, and bundled product recommendation. However, this task remains highly challenging: (1) product identifiers such as barcodes are often missing or inconsistently used, and (2) product names across platforms vary widely in structure, terminology, and granularity. Even small differences in attributes such as brand, variant, specification, or quantity can change whether two listings refer to identical products. This ambiguity makes reliable SKU mapping one of the most pressing technical bottlenecks in modern e-commerce pipelines [1, 2, 3].
Existing approaches to SKU mapping largely fall into rule-based systems and keyword-driven similarity methods. Rule-based systems rely on explicit heuristics, such as string normalization, brand token matching, or quantity checks. While effective in narrow cases, these methods struggle to generalize across product categories and cannot capture nuanced distinctions [4]. Keyword similarity methods, on the other hand, often overemphasize surface overlaps while ignoring semantic differences in brand origin, bundle configuration, or specifications. As a result, both approaches frequently yield false positives or negatives, particularly in ambiguous product families and bundled offerings [5].
Instead of pursuing model fine-tuning as existing studies suggest [6, 7, 8, 9], we take a different direction: Can a new SKU mapping framework be built directly on top of general-purpose LLMs, without costly and inflexible retraining? Continuous fine-tuning is not only expensive but hard to maintain under the dynamic nature of e-commerce, where product catalogs evolve rapidly and diverge from training distributions. Moreover, SKU mapping inherently requires access to up-to-date product information, making web search indispensable: attributes such as availability, branding, or origin change frequently, and static models cannot reliably capture these shifts. Our goal is to treat the LLM as a flexible reasoning module that can adapt immediately to new products, remain cost-efficient, and avoid the overhead of retraining with each update.
To address these challenges, we propose Question-to-Knowledge (Q2K), a novel multi-agent framework that transforms SKU mapping into a process of generating and validating inspectable facts. Compared to prior methods that rely solely on surface similarity or opaque model predictions, Q2K explicitly decomposes the mapping task into three coordinated agents:
Reasoning Agent
which analyzes a base–candidate product pair and generates targeted attribute-specific disambiguation questions (e.g., brand, origin, specification, quantity) to identify missing or ambiguous features;
Knowledge Agent
which conducts focused web searches to resolve these questions by synthesizing authoritative evidence into concise, self-contained answers;
Deduplication Agent
which leverages dense retrieval to detect and reuse previously stored question–answer reasoning traces, reducing redundancy, lowering cost, and ensuring consistency across mappings.
Furthermore, a human-in-the-loop mechanism enhances Q2K by validating low-confidence cases, with corrections fed back into the reasoning database for continuous improvement. This design allows Q2K to produce transparent, reusable, and cost-efficient reasoning chains rather than opaque, one-off judgments.
To assess the effectiveness of Q2K, we evaluate it against $4$ representative baselines: (1) a rule-based mapping system combining string normalization, brand alignment, and quantity matching heuristics; (2) zero-shot inference, where models directly judge equivalence from product names; (3) few-shot inference, which augments prompts with a small set of labeled examples; (4) web-search inference, where models are allowed to query the web without structured decomposition.
Across real-world consumer goods datasets, Q2K consistently outperforms strong baselines, delivering significant improvements in SKU mapping accuracy. Crucially, by retrieving top-$k$ relevant reasoning traces rather than repeatedly issuing new web searches, Q2K lowers computational and operational costs while preserving robustness in challenging cases such as bundle identification, brand–origin disambiguation, and multi-source sourcing. These results highlight Q2K’s ability to tackle one of the most persistent challenges in e-commerce. By reframing SKU mapping as an interpretable, retrieval-augmented reasoning process, Q2K combines automation with human oversight to enable scalable and trustworthy product integration across heterogeneous platforms.
II Data Collection
We constructed a dataset of product name pairs through a semi-manual collection process, designed in close collaboration with our industry partner. The collection procedure involved specialized product sellers as human annotators, who possess domain expertise in distinguishing subtle variations across brands, specifications, and bundled offerings. To further ensure data quality, their annotations were regularly cross-validated through supervision checks conducted by senior reviewers.
II-A Collection Procedure
The client initially provided a list of reference product names, denoted as base_product. Human data collectors were then instructed to search a designated e-commerce website and gather a list of potentially matching product names. For each candidate, the collector extracted the product title and recorded it as compared_product.
In addition to recording potential matches, the collectors explicitly flagged products that did not correspond to the given base_product. These flagged items served as negative examples for subsequent product mapping. To further ensure accuracy, a supervisor reviewed the flagged items and validated whether they were labeled correctly.
II-B Preprocessing
Minimal preprocessing was applied to maintain fidelity to the raw data. Each instance was represented as a pair:
$(\textit{base\_product},\textit{compared\_product}).$
. No additional normalization (e.g., tokenization, stopword removal, or lemmatization) was performed at this stage in order to preserve the original wording of product listings.
II-C Dataset Statistics
The resulting dataset contains approximately $72,250$ rows of product pairs. Notably, the majority of collected samples fall within the food & beverage (F&B) domain. Based on manual inspection, we estimate that around $3\%$$5\%$ of the pairs may contain labeling errors, largely due to subtle attribute differences that were difficult for annotators to resolve.
Base Product Compared Product
MegaDoseD Vitamin D3 4000IU, 120 tablets × 3 (12 months) Korea Eundan MegaDoseD Vitamin D3 4000IU Swiss-made, 120 tablets × 3 (12 months)
TABLE I: Example of a mislabeled pair. Despite sharing identical dosage and packaging, the pair was annotated as non-equivalent since the compared product emphasized its Swiss origin. In reality, both refer to the same product line, highlighting the difficulty of annotation.
This example illustrates the reason why SKU mapping is inherently challenging: optional descriptors such as origin, supplier emphasis, or promotional claims may appear decisive but do not always indicate genuine product differences. Despite such cases, the dataset remains a representative benchmark for evaluating the proposed framework in realistic e-commerce settings.
Figure 1: Overview of the Q2K workflow. The Reasoning Agent generates disambiguation questions from a product-pair, the Deduplication Agent retrieves top-k reasoning traces and checks information gain, and the Knowledge Agent provides authoritative answers when evidence is insufficient.
III Methodology
Our proposed framework Q2K operates through the collaboration of $3$ specialized agents: (1) a Reasoning Agent, (2) a Knowledge Agent, and (3) a Deduplication Agent. An overview of the framework is illustrated in Figure 1.
III-A Reasoning Agent
Product mapping often fails due to missing or ambiguous attribute information, such as unclear manufacturer origin, inconsistent quantity representation, or variant specifications. Motivated by this observation, the Reasoning Agent decomposes each product-pair comparison into targeted, attribute-specific questions. Given a base product name $P_{b}$ and a candidate product name $P_{c}$, the agent generates a set of disambiguation questions:
$\mathcal{Q}=\{q_{1},q_{2},\dots,q_{m}\}=G_{\mathcal{R}}(P_{b},P_{c}),$ (1)
where $G_{\mathcal{R}}$ denotes the Reasoning Agent, and $m$ is dynamically determined based on the ambiguity detected in the product pair. Each generated question explicitly corresponds to one of five matching dimensions: brand, core product name, variant, specification, or quantity (see Table II).
Attribute Role in Question Generation
Brand Ensures consistency of manufacturer, label, or maker across both products.
Core Product Name Identifies whether names are synonyms or distinct product lines.
Variant Clarifies options such as flavor, color, or finish.
Specification Validates differences in size, weight, volume, or technical details.
Quantity Confirms pack size, unit count, or bundle configuration.
TABLE II: Five matching dimensions used by the Reasoning Agent to generate targeted disambiguation questions for SKU mapping.
III-B Knowledge Agent
While the Reasoning Agent formulates targeted questions, the Knowledge Agent retrieves factual evidence to resolve them. For each generated question $q_{i}\in\mathcal{Q}$, the Knowledge Agent executes focused web queries and synthesizes authoritative answers into structured outputs:
$a_{i}=G_{\mathcal{K}}(q_{i},P_{b},P_{c}),$ (2)
where $G_{\mathcal{K}}$ denotes the Knowledge Agent, and $a_{i}$ is a concise, self-contained explanation oriented toward product equivalence checking. All answers are compiled into an evidence set:
$\mathcal{A}=\{a_{1},a_{2},\dots,a_{m}\}.$ (3)
This process ensures that ambiguous attributes are resolved with external evidence, rather than relying solely on LLM priors. By grounding reasoning in inspectable facts, Q2K supports both automation and human validation.
III-C Deduplication Agent
Naively executing web searches for every product-pair comparison introduces redundancy and high operational cost. To address this, the Deduplication Agent coordinates two key operations: (1) retrieving potentially relevant reasoning traces, and (2) deciding whether these traces provide sufficient information to resolve the current case.
First, instead of treating each question independently, the set of generated questions $\{q_{1},q_{2},\dots,q_{m}\}$ is concatenated into a structured sequence:
$Q_{\text{concat}}=\text{``Question1: }q_{1}\text{;}\dots\text{; Question}m:q_{m}\text{''}.$
This concatenated query string is embedded using a sentence transformer and compared against the repository of previously stored reasoning traces. The top-$k$ most semantically similar entries are retrieved:
$R_{k}(Q_{\text{concat}})=\operatorname{Top-}k\left(\cos(E(Q_{\text{concat}}),E(\mathcal{Q}_{\text{DB}}))\right),$ (4)
where $E(\cdot)$ is the embedding function and $\mathcal{Q}_{\text{DB}}$ denotes the repository of concatenated question–answer reasoning traces. Formally, we define
$\mathcal{Q}_{\text{DB}}=\{(Q_{1},A_{1}),(Q_{2},A_{2}),\dots,(Q_{j},A_{j})\},$
where each $Q_{i}$ is a concatenated sequence of disambiguation questions and $A_{i}$ is its corresponding validated answer set. Thus, $\mathcal{Q}_{\text{DB}}$ represents $j$ pre-saved knowledge pairs that can be directly reused to resolve new product-pair comparisons.
The Deduplication Agent subsequently evaluates whether the retrieved traces provide sufficient information gain to support equivalence reasoning. If the retrieved evidence is adequate, the stored answers are directly reused, ensuring both efficiency and consistency. If the evidence is insufficient, the agent escalates to the full Reasoning–Knowledge process: new disambiguation questions are issued, authoritative answers are synthesized via web search, and the resulting reasoning chain is added back into the knowledge base.
IV Experiments
IV-A Setup
We conduct experiments on our collected SKU mapping dataset to evaluate the effectiveness of the proposed Question-to-Knowledge (Q2K) framework. All experiments are designed to ensure fairness in comparison with baseline methods by using the same product-pair inputs and evaluation protocol.
IV-A1 LLM and Retrieval Models
We use GPT-4o [10] for all three agents in Q2K, accessed via API with temperature fixed at 0 to ensure deterministic and reproducible outputs.
For retrieval, we employ text-embedding-3-small model from OpenAI [10] to embed concatenated question sequences. The top-$5$ most similar reasoning traces are retrieved from the knowledge repository for ensuring a balance between evidence coverage and efficiency.
IV-A2 Evaluation Metrics
Since the SKU mapping task is framed as a binary classification problem (equivalent vs. non-equivalent), we report accuracy as the primary evaluation metric. Formally, given $N$ product pairs with ground-truth labels $y_{i}\in\{0,1\}$ and predicted labels $\hat{y}_{i}$, accuracy is defined as:
$\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\{y_{i}=\hat{y}_{i}\},$ (5)
where $\mathbf{1}\{\cdot\}$ is the indicator function. Accuracy directly reflects the proportion of correctly mapped product pairs.
IV-A3 Baselines
We compare Q2K against $4$ baselines:
  • Rule-Based Matching: This baseline implements a deterministic pipeline built on handcrafted heuristics commonly used in e-commerce product matching. The procedure consists of three sequential stages:
    1. 1.
      String Normalization: Product titles are lowercased and stripped of punctuation, stopwords, and special characters. Common units and symbols are standardized to canonical forms.
    2. 2.
      Brand Token Matching: A curated dictionary of brand names is used to detect and align brand tokens in both the base and compared product names. If brand tokens are mismatched or absent, the pair is labeled as non-equivalent.
    3. 3.
      Quantity and Specification Alignment: Numerical tokens are extracted using regular expressions and compared across products. Bundle-related attributes must align for equivalence.
    The final decision rule is: two products are considered equivalent only if their normalized core tokens, detected brand, and extracted numerical specifications all match.
  • Zero-Shot Inference: GPT-4o directly predicts product equivalence from raw product names without additional context. This tests the raw reasoning ability of LLMs.
  • Few-Shot Inference: GPT-4o is prompted with a small set of labeled examples before inference. This aims to capture attribute diversity through in-context learning.
  • Web-Search Inference: GPT-4o is equipped with unrestricted web search for each product pair, without structured decomposition. This allows factual grounding but incurs higher latency and cost.
IV-B Experiment Results
Our experiment results are shown in table III. Zero-shot prompting directly compares product names without auxiliary context. While computationally efficient, this approach often struggles with ambiguous attributes such as bundle size or variant specification, leading to misclassifications. It achieves an accuracy of $90.34\%$, showing that LLMs can capture some semantic overlap but remain limited by under-specified product details. Few-shot prompting introduces exemplar cases to guide inference and improves performance to $92.68\%$, yet its reliance on a fixed set of examples reduces adaptability to novel product formats. Incorporating web search further enhances grounding by retrieving external evidence, raising accuracy to $93.44\%$, though this comes at the cost of additional latency.
Method Accuracy (Binary Classification)
Rule-Based Matching 0.6737
Zero-Shot Inference 0.9034
Few-Shot Inference 0.9268
Web Search Inference 0.9344
Query2Knowledge (Q2K) 0.9562
TABLE III: Comparison of SKU mapping accuracy across baseline methods and our proposed framework. Q2K (Ours) achieves the highest accuracy by combining query decomposition with web-based evidence retrieval.
In contrast, our proposed framework Q2K leverages query decomposition and multi-agents to systematically resolve ambiguous attributes with reusable reasoning chains. By combining structured disambiguation with web-based evidence retrieval, Q2K achieves the best performance with $95.62\%$ accuracy. This demonstrates its effectiveness in addressing the key limitations of prior baselines: limited adaptability in zero/few-shot methods and one-off reasoning in web-search prompting. The results highlight Q2K’s ability to provide scalable automation and interpretable, high-fidelity SKU mapping.
IV-C Ablation Study
To better understand the contribution of each component in Q2K, we analyze the internal dynamics of the framework. On average, the Reasoning Agent generated $1.4$ attribute-specific disambiguation questions per product-pair comparison, reflecting that most cases required clarification of a single ambiguous attribute, while a minority involved multiple attributes.
Ablation Metric Value
Average questions generated per product pair $1.4$
Cases requiring multiple questions Minority ( $20\%$ )
Deduplication retrieval activation rate $22\%$
TABLE IV: Ablation study of Q2K components. The Reasoning Agent typically generates a small number of questions, while the Deduplication Agent reuses prior reasoning in $22\%$ of cases, reducing cost and redundancy.
We further examined the role of the Deduplication Agent in reusing previously validated reasoning traces. Retrieval was activated in approximately $22\%$ of cases, where the system successfully leveraged existing question–answer pairs instead of performing new web searches. This significantly reduced redundancy and lowered the number of costly web queries, while maintaining consistent judgments across similar products.
V Conclusion
We presented Question-to-Knowledge (Q2K), a multi-agent framework for SKU mapping that integrates query decomposition, web-based evidence retrieval, and reasoning trace reuse. Experiments on real-world consumer goods datasets show that Q2K achieves $95.62\%$ accuracy, surpassing rule-based, zero-shot, few-shot, and web-search baselines while reducing redundant computation through selective retrieval. By grounding decisions in inspectable facts and maintaining a reusable knowledge base, Q2K provides interpretable, scalable, and cost-efficient SKU mapping. These results highlight the potential of multi-agent reasoning frameworks to advance reliable product integration in dynamic e-commerce environments.